# Dynamic annihilation pathways of magnetic skyrmions

### v1

@@ Acknowledgments
Author Contributions
M.C. and E.I. conceived and developed the project. M.C. implemented the 2D pseudospectral model with Dzyaloshinskii-Moriya interaction. Both authors wrote the manuscript.
Corresponding authors
Correspondence to
eiacocca@uccs.edu
mcopus@uccs.edu
Supplementary material: Understanding the physical limits of topological protection via the pseudospectral annihilation of magnetic skyrmions
Matthew Copus and Ezio Iacocca
SI 1 Derivation of 2D Dzyaloshinskii-Moriya Interaction PS-LL Kernel
Figure SI 1: Geometry to derive the DMI based on an element at position $\{i,j\}$. The distances are represented by the vector r.
Consider a 2D ferromagnet in the $\hat{x}\hat{z}$-plane that is subject an interfacial Dzyaloshinskii–Moriya interaction (DMI) with the symmetry breaking occurring in the $\hat{z}$ direction. The DMI energy for a site $\mathbf{S}_{(i,j)}$ due to its four nearest neighbour interactions is given by
$U_{DMI}$ $=$ $\mathbf{D}_{(i-1,i)}\cdot\left(\mathbf{S}_{(i-1,j)}\times\mathbf{S}_{(i,j)}\right)+\mathbf{D}_{(i,i+1)}\cdot\left(\mathbf{S}_{(i,j)}\times\mathbf{S}_{(i+1,j)}\right)+$ (SI 1)
$\mathbf{D}_{(j-1,j)}\cdot\left(\mathbf{S}_{(i,j-1)}\times\mathbf{S}_{(i,j)}\right)+\mathbf{D}_{(j,j+1)}\cdot\left(\mathbf{S}_{(i,j)}\times\mathbf{S}_{(i,j+1)}\right)$
@@
The DMI vectors are given by the cross product between the direction of the vector $\mathbf{r}$ connecting two points (shown in figure 1) and the symmetry breaking direction
$\mathbf{D}_{(i-1,i)}$ $=$ $\mathbf{D}_{(i,i+1)}=D(\hat{r}_{(i-1,i)}\times\hat{z})=D(\hat{x}\times\hat{z})=-D\hat{y},$ (SI 3a)
$\mathbf{D}_{(j-1,j)}$ $=$ $\mathbf{D}_{(j,j+1)}=D(\hat{r}_{(j-1,j)}\times\hat{z})=D(\hat{y}\times\hat{z})=D\hat{x}.$ (SI 3b)
@@
$\mathbf{H}_{DMI}=-\frac{\partial U_{DMI}}{\partial S_{x(i,j)}}\hat{x}-\frac{\partial U_{DMI}}{\partial S_{y(i,j)}}\hat{y}-\frac{\partial U_{DMI}}{\partial S_{z(i,j)}}\hat{z}.$ (SI 5)
After taking the derivatives, we get the effective field acting on site $\mathbf{S}_{(i,j)}$ due to DMI from its nearest neighbours
$H_{x}$ $=$ $D\left(S_{z(i-1,j)}-S_{z(i+1,j)}\right),$ (SI 6a)
$H_{y}$ $=$ $D\left(S_{z(i,j-1)}-S_{z(i,j+1)}\right),$ (SI 6b)
$H_{z}$ $=$ $D\left(S_{x(i+1,j)}-S_{x(i-1,j)}+S_{y(i,j+1)}-S_{y(i,j-1)}\right).$ (SI 6c)
We now make the assumption that site $\mathbf{S}_{(i,j)}$ undergoes wave motion and that the neighbouring sites experience the same motion with a phase shift
$S_{z(i-1,j)}$ $=$ $S_{z(i,j)}e^{-i(k_{x}a)},$ (SI 7a)
$S_{z(i+1,j)}$ $=$ $S_{z(i,j)}e^{i(k_{x}a)},$ (SI 7b)
$S_{z(i,j-1)}$ $=$ $S_{z(i,j)}e^{-i(k_{y}a)},$ (SI 7c)
$S_{z(i,j+1)}$ $=$ $S_{z(i,j)}e^{i(k_{y}a)},$ (SI 7d)
$S_{x(i+1,j)}$ $=$ $S_{x(i,j)}e^{i(k_{x}a)},$ (SI 7e)
$S_{x(i-1,j)}$ $=$ $S_{x(i,j)}e^{-i(k_{x}a)},$ (SI 7f)
$S_{y(i,j+1)}$ $=$ $S_{y(i,j)}e^{i(k_{y}a)},$ (SI 7g)
$S_{y(i,j-1)}$ $=$ $S_{y(i,j)}e^{-i(k_{y}a)}.$ (SI 7h)
@@
$H_{x}$ $=$ $D\left(S_{z(i,j)}e^{-i(k_{x}a)}-S_{z(i,j)}e^{i(k_{x}a)}\right),$ (SI 8a)
$H_{y}$ $=$ $D\left(S_{z(i,j)}e^{-i(k_{y}a)}-S_{z(i,j)}e^{i(k_{y}a)}\right),$ (SI 8b)
$H_{z}$ $=$ $D\left(S_{x(i,j)}e^{i(k_{x}a)}-S_{x(i,j)}e^{-i(k_{x}a)}+S_{y(i,j)}e^{i(k_{y}a)}-S_{y(i,j)}e^{-i(k_{y}a)}\right).$ (SI 8c)
@@
$dx:$ $\mathcal{F}^{-1}\left[-\hat{S}_{z(i,j)}D2i\sin{(k_{x}a)}\right],$ (SI 10a)
$dy:$ $\mathcal{F}^{-1}\left[-\hat{S}_{z(i,j)}D2i\sin{(k_{y}a)}\right],$ (SI 10b)
$dz:$ $\mathcal{F}^{-1}\left[\hat{S}_{x(i,j)}D2i\sin{(k_{x}a)}+\hat{S}_{y(i,j)}D2i\sin{(k_{y}a)}\right].$ (SI 10c)

### v3

@@ Acknowledgments
Data Availability
The data that support the findings of this article are openly available [62].
Appendix A Derivation of the 2D Dzyaloshinskii-Moriya Interaction PS-LL Kernel
Figure 6: Geometry to derive the DMI based on an element at position $\{i,j\}$. The distances are represented by the vector r.
Consider a 2D ferromagnet in the $\hat{x}\hat{z}-$plane that is subject to an interfacial Dzyaloshinskii–Moriya interaction (DMI) with the symmetry breaking occurring in the $\hat{z}$ direction. The DMI energy for a site $\mathbf{S}_{(i,j)}$ due to its four nearest neighbor interactions is given by
$U_{DMI}$ $=$ $\mathbf{D}_{(i-1,i)}\cdot\left(\mathbf{S}_{(i-1,j)}\times\mathbf{S}_{(i,j)}\right)+$ (15)
$\mathbf{D}_{(i,i+1)}\cdot\left(\mathbf{S}_{(i,j)}\times\mathbf{S}_{(i+1,j)}\right)+$
$\mathbf{D}_{(j-1,j)}\cdot\left(\mathbf{S}_{(i,j-1)}\times\mathbf{S}_{(i,j)}\right)+$
$\mathbf{D}_{(j,j+1)}\cdot\left(\mathbf{S}_{(i,j)}\times\mathbf{S}_{(i,j+1)}\right)$
@@ Appendix A Derivation of the 2D Dzyaloshinskii-Moriya Interaction PS-LL Kernel
The DMI vectors are given by the cross product between the direction of the vector $\mathbf{r}$ connecting two points (shown in Fig. 6) and the symmetry breaking direction
[…]
$=$ $D(\hat{x}\times\hat{z})=-D\hat{y}$,
[…]
$=$ $D(\hat{y}\times\hat{z})=D\hat{x}$.
@@
$\mathbf{H}_{DMI}=-\frac{\partial U_{DMI}}{\partial S_{x(i,j)}}\hat{x}-\frac{\partial U_{DMI}}{\partial S_{y(i,j)}}\hat{y}-\frac{\partial U_{DMI}}{\partial S_{z(i,j)}}\hat{z}$. (19)
After taking the derivatives, we get the effective field acting on site $\mathbf{S}_{(i,j)}$ due to DMI from its nearest neighbors
$H_{x}$ $=$ $D\left(S_{z(i-1,j)}-S_{z(i+1,j)}\right),$ (20a)
$H_{y}$ $=$ $D\left(S_{z(i,j-1)}-S_{z(i,j+1)}\right),$ (20b)
$H_{z}$ $=$ $D\Big(S_{x(i+1,j)}-S_{x(i-1,j)}$ (20c)
$+S_{y(i,j+1)}-S_{y(i,j-1)}\Big).$
We now make the assumption that site $\mathbf{S}_{(i,j)}$ undergoes wave motion and that the neighboring sites experience the same motion with a phase shift
$S_{z(i-1,j)}$ $=$ $S_{z(i,j)}e^{-i(k_{x}a)},$ (21a)
$S_{z(i+1,j)}$ $=$ $S_{z(i,j)}e^{i(k_{x}a)},$ (21b)
$S_{z(i,j-1)}$ $=$ $S_{z(i,j)}e^{-i(k_{y}a)},$ (21c)
$S_{z(i,j+1)}$ $=$ $S_{z(i,j)}e^{i(k_{y}a)},$ (21d)
$S_{x(i+1,j)}$ $=$ $S_{x(i,j)}e^{i(k_{x}a)},$ (21e)
$S_{x(i-1,j)}$ $=$ $S_{x(i,j)}e^{-i(k_{x}a)},$ (21f)
$S_{y(i,j+1)}$ $=$ $S_{y(i,j)}e^{i(k_{y}a)},$ (21g)
$S_{y(i,j-1)}$ $=$ $S_{y(i,j)}e^{-i(k_{y}a)}.$ (21h)
@@
$H_{x}$ $=$ $D\left(S_{z(i,j)}e^{-i(k_{x}a)}-S_{z(i,j)}e^{i(k_{x}a)}\right),$ (22a)
$H_{y}$ $=$ $D\left(S_{z(i,j)}e^{-i(k_{y}a)}-S_{z(i,j)}e^{i(k_{y}a)}\right),$ (22b)
$H_{z}$ $=$ $D\Big(S_{x(i,j)}e^{i(k_{x}a)}-S_{x(i,j)}e^{-i(k_{x}a)}$ (22c)
$+S_{y(i,j)}e^{i(k_{y}a)}-S_{y(i,j)}e^{-i(k_{y}a)}\Big).$
@@
$dx:$ $\mathcal{F}^{-1}\left[-\hat{S}_{z(i,j)}D2i\sin{(k_{x}a)}\right],$ (24a)
$dy:$ $\mathcal{F}^{-1}\left[-\hat{S}_{z(i,j)}D2i\sin{(k_{y}a)}\right],$ (24b)
$dz:$ $\mathcal{F}^{-1}\Big[\hat{S}_{x(i,j)}D2i\sin{(k_{x}a)}$ (24c)
$+\hat{S}_{y(i,j)}D2i\sin{(k_{y}a)}\Big].$